OpenStudy (zyberg):

Is a^2 + b^2 coprime to a and b, if a and b are coprime between each other?

ganeshie8 (ganeshie8):

Hi

OpenStudy (zyberg):

Hello there, @geneshie8

OpenStudy (zyberg):

Still struggling quite hard with coprimes :/

ganeshie8 (ganeshie8):

Let's try proof by contradiction

ganeshie8 (ganeshie8):

Suppose p is a prime factor that divides both a and a^2 + b^2

ganeshie8 (ganeshie8):

Then there exist integers a' and c' that satisfy a = pa' a^2 + b^2 = pc'

ganeshie8 (ganeshie8):

With me so far ?

jhonyy9 (jhonyy9):

hi everybody ! @ganeshie8 i think from my past 4 years experience from here and different math. forums so i saw that peopels dont like - or dont know it how - working with proof by contradiction method - i dont know it way bc. this is easy - my opinion

jhonyy9 (jhonyy9):

sorry - i dont know it why ?

ganeshie8 (ganeshie8):

Yeah I see contradiction proofs very rarely here. But I do know a few ppl here who love contradiction so much that they always try it as their first attempt. I'm one of them haha

OpenStudy (kainui):

That was much shorter than my way, I am not even sure my way works 100% there are a few details wrong I think with my path.

OpenStudy (kainui):

Let's see here, we know from the problem that \(\gcd(a,b)=1\) and we want to see if these statements are true or not, \(\gcd(a^2+b^2,a)=1\) and \(\gcd(a^2+b^2,b)=1\). I decided to just pick one and go with it. \[\gcd(a^2+b^2,a)=\gcd(a^2-2ab+b^2,a)\] So like you told us the other day, kY-kX = k(Y-X). We could repeat this multiple times though, kY-kX-kX = k(Y-2X). So I decided to subtract of 2b times of a from \(a^2+b^2\) to get \(a^2-2ba+b^2\) above, this is all I've done in that step. \[\gcd(a^2-2ab+b^2,a) = \gcd((a+b)(a-b),a)\] Next step I took was to factor it (that was my motivation in case you were wondering.) \[\gcd((a+b)(a-b),a) \le \gcd(a+b,a)*\gcd(a-b,a)\] Now let's check this out, probably best seen with an example, \(\gcd(8,2)\le \gcd(2,2)*\gcd(4,2)\) which is the same as saying \(2 \le 4\). Basically by separating it apart we get this sort of inequality. This seems spooky, and this is definitely a brave move since inequalities might not help. \[\gcd(a+b,a)=1\]\[\gcd(a-b,a)=1\] We know these are true from the reasoning with kY-kX=k(Y-X) from earlier. So hey, if you string it altogether we get: \[\gcd(a^2+b^2) \le 1\] Which, since the gcd function can't take any values less than 1, it ends up being equal to 1! I think we can repeat a similar process for b, although maybe not, since I think perhaps the a-b vs b-a term ends up somehow not working out exactly the same. This is just what I did, and I hope it helps but it probably isn't the only way.

ganeshie8 (ganeshie8):

I think it works nicely I notice a small error a^2 - 2ab + b^2 is not (a +b)(a-b) but that doest make the proof wrong, the overall structure should work I guess

OpenStudy (zyberg):

@Kainui why did you subtract 2ab from left number, but not from right? I kind of understand why you would subtract a, but where does 2b come from?

OpenStudy (zyberg):

Or, I get where from is 2a, but I don't get where is b from.

OpenStudy (kainui):

Haha wow yeah, I like don't even know how to multiply I guess anymore. \(a^2-2ab+b^2 = (a-b)^2\). b is just some number, and it comes from the problem itself.

OpenStudy (zyberg):

Oh, so we can subtract anything we want, not only one number from another?

OpenStudy (kainui):

b is a number

OpenStudy (zyberg):

Yes. But why can we subtract it from a^2 + b^2? Is it because GCD(a, b)= 1, so a and b won't have common divisors and that leads to it?

OpenStudy (kainui):

I know it's kinda weird so here, I'll sorta write it in steps of taking away \(a\), \[\gcd(a^2+b^2,a)\]\[ =\gcd(a^2+b^2-a,a) \]\[=\gcd(a^2+b^2-2*a,a) \]\[=\gcd(a^2+b^2-3*a,a) \]\[=\cdots \]\[=\gcd(a^2+b^2-(2b-1)*a,a)\]\[=\gcd(a^2+b^2-2b*a,a)\]

OpenStudy (kainui):

we could take \(a\) away 3 times, 17 times, whatever. Now let's replace and let's say \(x=3\) and \(y=17\). We can take away \(a\) for \(x\) times too. The main thing is, we know we aren't taking away more \(a\)s than we have left of \(a^2+b^2\). Fortunately we know that \(a^2-2ab+b^2 = (a-b)^2 \ge 0\) so we're allowed cause anything squared is greater than or equal to zero.

OpenStudy (zyberg):

Okay, I think that I understood it now. :) Another question: why did you say that GCD(a - b, a) = 1? Is it from the same thing like you wrote there? But GCD(a, a)= a, so how would you subtract b from it, if it's not really possible?

OpenStudy (kainui):

Also, try plugging in some values of \(a\) and \(b\) if you get confused, just make sure whatever you plug in obeys \(\gcd(a,b)=1\). Making up your own concrete examples is always good.

ganeshie8 (ganeshie8):

Remember that logs example ? How gcd can be obtained by taking the difference of the logs...

OpenStudy (kainui):

Yeah so this last bit I was trying to be tricky, and it may or may not work. Luckily we don't have to, since @ganeshie8 pointed out my mistake we actually don't have to worry about this at all. But suppose you still want to worry about it anyways, just for fun. Then think about it :P

OpenStudy (kainui):

I say it won't work because I am specifically forced to assume a > b due to my faulty reasoning in factoring earlier btw

OpenStudy (zyberg):

@geneshie8 , yes :) But is it valid to take away a log with length of b from already cut logs into a length pieces? Looking at GCD(a - b, a) ;) I guess that it would be possible to GCD(b, a) *(-1) = GCD(-b, -a) but then it makes even less sense :/ Oh, that clears up everything, kanui ;)

ganeshie8 (ganeshie8):

Try this later maybe Suppose the logs are of lengths 35 and 14 How are you gona cut them into equal pieces such that the length of the piece is maximum ?

OpenStudy (zyberg):

@ganeshie8 yes, gcd would be 7. 35 - 14 = 21 -14 = 7; 14 - 7 = 7; 1st and 2nd numbers of gcd are equal, gcd is equal to 7.

OpenStudy (zyberg):

Thank you for your help, everyone! :)

ganeshie8 (ganeshie8):

gcd(35, 14) = gcd(35-2*14, 14)

ganeshie8 (ganeshie8):

You can subtract any multiple of the other integer. In general \[\gcd(a,b)= \gcd(a-nb, b)\]

OpenStudy (zyberg):

Yes, I didn't notice that b is a multiple of a this time. :)

ganeshie8 (ganeshie8):

I challenge you to prove above fact. You have all the tools necessary to prove by now

ganeshie8 (ganeshie8):

Try those proofs when free I'm sure you will find them fun after doing a few

OpenStudy (zyberg):

I accept the challenge! :) if a > b gcd(a, b) = gcd(a - nb, b) gcd(a, b) = p a = pa` b = pb` a - b = pa` - pb` = p(a` - b`) = c if c > b gcd(c, b) = gcd(c - b, b) c - b = p(a` - b`) - pb` = p(a` - 2b`) = d And if d > b it can be repeated. So, for a = nb (or a` = nb`) it works to subtract b n times from a. Is it a good proof?

ganeshie8 (ganeshie8):

Brilliant! But why are you restricting the difference to be positive ? Are you scared of negative numbers ?

OpenStudy (zyberg):

Shouldn't gcd function work with only positive values? ;)

ganeshie8 (ganeshie8):

gcd(a, b) = gcd(a-nb) This is true for all integers n, including negatives

OpenStudy (zyberg):

Oh, okay. That is addition, in other words. Thanks for pointing it out! :)

ganeshie8 (ganeshie8):

In our earlier example gcd(35, 14) = gcd(35-100*14, 14)

ganeshie8 (ganeshie8):

I'd like you to verify above

OpenStudy (zyberg):

Well, it works, since gcd(a, b)=gcd(a - nb, b) And to check if it really works, we can look if gcd(35, 14) = gcd(-1365, 14) sine gcd(35, 14) = 7, we try to divide -1365 by 7 and we get -195, so it is right! (we know that it is the biggest common divisor and there can't be bigger ones because 14 wouldn't divide from 7 < numbers)

ganeshie8 (ganeshie8):

Ofcourse you can't have negative length logs. gcd can never be negative because any negative integer is always less than 1. And what's special about 1 ?

ganeshie8 (ganeshie8):

Looks good, so gcd can be defined for negative integers too even though gcd is never negative

OpenStudy (zyberg):

Oh, it was still about logs :D 1 is the least natural divisor. That is why it is special. gcd(a,b) = 1 would mean that a and b are coprime.

ganeshie8 (ganeshie8):

Yes. More importantly 1 is a divisor of every integer.

ganeshie8 (ganeshie8):

So there is a natural lower limit for the set of gcds.

OpenStudy (zyberg):

I think that slowly I am getting to the point where I start to understand gcd and coprimes a lot better, thanks to you ;)

ganeshie8 (ganeshie8):

Np. You're doing great. Keep up the good work :)

ganeshie8 (ganeshie8):

Hey did you finish the contradiction proof that I've started earlier ?

ganeshie8 (ganeshie8):

I cannot type long replies as I'm on mobile with a cheap touch kb

OpenStudy (zyberg):

Nope. I can't see where to go from the step you wrote :/ Should I try a^2 + b^2 - a some times to get p?

ganeshie8 (ganeshie8):

No You see two equations in my reply there Simply substitute first equation into second

ganeshie8 (ganeshie8):

Then stare at the resulting equation You will see what to do next hopefully

ganeshie8 (ganeshie8):

a = pa' a^2 + b^2 = pc' Replacing a by pa' gives (pa')^2 + b^2 = pc' rearranging gives b^2 = p(some integer) That means p divides b. Contradiction.

OpenStudy (zyberg):

Tried to do something similar, but didn't notice that it's contradiction. But since a and b are coprime wouldnt b = pb` as well? So where does the contradiction come from? @ganeshie8

OpenStudy (zyberg):

Oh, from where a and b are coprime and the p would be 1?

ganeshie8 (ganeshie8):

We have started with a = pa' Tell me how can p divide b too ?

ganeshie8 (ganeshie8):

We have assumed p is prime. p is not 1.

ganeshie8 (ganeshie8):

That is where the contradiction cones from

OpenStudy (zyberg):

But well, a and b are coprime... Ohh! Yes, we assumed that a isn't coprime with a^2 + b^2! That's amazing! Your way is quite simple, actually :O Just needed to notice contradiction... thank you very much!

ganeshie8 (ganeshie8):

Haha yeah contradiction proofs are always elegant. You may lookup Euclid's proof of infinitude of primes for more motivation :)

OpenStudy (zyberg):

Mm, that sounds interesting :) Going to check it out this evening (I noticed that Euclid comes up a lot when talking about primes).

ganeshie8 (ganeshie8):

Have fun !